OpenStudy (anonymous):

Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (anonymous):

4x-7=20 +7=+7 4x=27 /4=/4

OpenStudy (anonymous):

and I know it's right because it is just right lol.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)